# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v2

@@ 5.2.2 Results
We study, in our predictive regression framework, a one-month-ahead regression $\text{Inflation}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{Unrate}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes all transformed covariates. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the unemployment rate in forecasting inflation. The predictive form is of particular interest among policymakers in leveraging the relationship as a practical tool.
@@
Table 4: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ : Full sample and subperiods
Sample Periods IVX Dlasso XDlasso
Est. S.E. Est. S.E. Est. S.E.
Full Sample (Jan. 1960 - Apr. 2024) -0.023** 0.010 0.019*** 0.006 -0.021 0.065
Pre-Volcker (Jan. 1960 - July 1979) 0.080 0.069 0.074*** 0.017 0.013 0.224
Volcker-Greenspan (Aug. 1979 - Jan. 2006) 0.035 0.063 -0.021 0.019 0.155 0.120
Bernanke-Yellen-Powell (Feb. 2006 - Apr. 2024) -0.044* 0.026 0.000 0.008 0.020 0.027
@@
Notes: ”Est.” and ”S.E.” represent the estimates and standard error, respectively. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. The tuning parameter for LASSO estimation is selected through 10-fold cross-validation. In XDlasso, instruments are generated based on (2.7) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
@@
Table 4 reports the point estimates and standard errors for $\theta_{1}^{*}$ using IVX, Dlasso and XDlasso across sample periods. The results by XDlasso in different subperiods are unanimous that the unemployment rate does not have significant predictive power in forecasting the inflation rate across all periods. These results add new insight to the recent debates on the Phillips curve: the predictive power of unemployment for inflation appears to be diminished when we control for other high dimensional time series as covariates. In contrast, IVX and Dlasso yield contradictory results. Dlasso delivers significantly positive estimates for the full sample and the Pre-Volcker period at the 1% significance level. Conversely, IVX detects significant negative effects for the full sample and Bernanke-Yellen-Powell period at the 5% and 10% significance levels, respectively. These discrepancies may be attributed to the failure of IVX to control for confounders and the neglect of the Stambaugh bias in Dlasso arising from nonstationarity. The empirical findings echo Mankiw, (2024)’s latest pessimistic remark: “The large confidence intervals for the natural rate, together with the apparent futility of this Holy Grail search, lead me to think that we should not expect much from the Phillips curve as a guide for forecasting inflation or for judging the stance of policy.”
@@ 6 Conclusion
One limitation of this paper is that it considers only the unit roots for persistent regressors. In addition to unit roots, there are many other models that portray the complex nonstationarity in real-world data, including local unit roots, fractional integration, cointegration, structural breaks, to name a few. LASSO’s consistency and inference under these time series models are yet to be studied, and one may attempt to generalize inference of the null to high dimensional parameters. These extension are currently being investigated.
@@ Appendix A Technical Proofs
[…]
Section A.1 proves Lemma 1 that establishes consistency of Slasso under general mixing conditions without an explicit linear process. Section 3.1 includes the proofs for the results in Section 3.1 that constructs consistency of the auxiliary LASSO regression adapted for bias correction. Section A.3 includes the proofs for the theorems in Section 3.2 about the asymptotic distribution of the XDlasso estimator and the order of its standard error, which determine the size and power of the XDlasso inference.
A.1 Proof of Lemma 1
The assumptions imposed in Section 3 slightly differ from those in Mei and Shi, (2024) (MS24, hereafter). Specifically, the linear process assumption in MS24 is extended to more general $\alpha$-mixing and sub-exponential conditions. The main modification of the proof is the Gaussian approximation error deduced in MS24’s Lemma B.4. The following lemma re-establishes Gaussian approximation errors for the stationary components $v_{t}$ defined as (3.1).
@@ A.1 Proof of Lemma 1
$\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(t\cdot V_{j}^{*})\right)\right|=O_{p}\left(\dfrac{(\log p)^{3/2}}{n^{1/4}}\right)$
@@
The convergence rate in Lemma A.1 is less sharp than that in MS24’s Lemma B.4, since we work with $\alpha$-mixing sequences.
The proof of Lemma 1 follows exactly the same roadmap of the proof of MS24’s Theorem 2. Main modifications lie in the convergence rates of the Gaussian approximation in the proof of MS24’s Proposition B.4. Specifically, we follow the notations in MS24 for this proof only. The convergence rate below MS24’s (B.48) becomes
$\sup_{j\in[p+1]}\sup_{0\leq t\leq n-1}n^{-1/2}|\xi_{jt}-\zeta_{jt}|=O_{p}\left(\dfrac{(\log p)^{3/2}}{n^{1/4}}\right).$
The last equality at Page 20 in MS24’s supplement becomes
$\|\widehat{\Sigma}-\widehat{\Upsilon}\|_{\infty}=O_{p}\left(n^{3/4}(\log p)^{3/2}\right).$
Then Equation (B.49) becomes
$\dfrac{\left|\delta^{\top}(\widehat{\Sigma}-\widehat{\Upsilon})\delta\right|}{n\|\delta\|_{2}^{2}}\leq 4L^{2}sO_{p}\left(n^{-1/4}(\log p)^{3/2}\right)=o_{p}\left(\dfrac{L^{-2}}{s\log p}\right)$
under Assumption 4 in the current paper, which implies $s^{2}=o(n^{1/4-\xi})$ for an arbitrarily small $\xi$.
We prove Lemma A.1 to complete assist the proof of Lemma 1.
@@
Recall that Assumption 2 imposes an upper bound for the $\alpha$-mixing coefficient. With $p=q=3$ in Equation (2.2) in the Corollary of Davydov, (1968, pp. 692), we have
@@
Step 2. Verifying (A.2). We use the strong Gaussian approximation from Lin and Lu, (1997)’s Theorem 9.3.1. Specifically, define $g(x)=\exp(x)$. By the sub-exponential tail imposed by Assumption 1, the sub-exponential norm of $\varepsilon_{k,t}$, denoted as $\|\varepsilon_{k,t}\|_{g}$ in Lin and Lu, (1997), is uniformly bounded by an absolute constant. It then suffices to verify the following two conditions required in the aforementioned theorem:
@@
    $\sum_{d=1}^{\infty}\alpha(d)^{1/4}\cdot\log(1/\alpha(d))<\infty$, where the parameter $\delta$ in Lin and Lu, (1997, Theorem 9.3.1) is taken as 2.
Then by Theorem 9.3.1 of Lin and Lu, (1997), for any $j\in[p]$
@@
Proof of (i). By (A.4),
$\dfrac{V_{k,t}}{t}=V_{k}^{*}-2\sum_{d=t+1}^{\infty}{\rm cov}_{k}(d)-\dfrac{2}{t}\sum_{d=1}^{t}d\cdot{\rm cov}_{k}(d)\to V_{k}^{*}$
as $t\to\infty$, where the limit applies (A.6) and (A.7). By Assumption 2, the long-run variance $V_{k}^{*}$ is bounded from below by some absolute constant. This result implies that $V_{k,t}/t$ is lower bounded by some absolute constant uniformly for all $(k,t)$.

### v3

@@ 5.2.2 Results
We study, in our predictive regression framework, a one-month-ahead regression $\text{Inflation}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{Unrate}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes high-dimensional covariates. Given the ambiguous persistence of the inflation rate discussed in the previous section, including the lagged dependent variable in our model may introduce further technical complications. Therefore, we do not recommend its inclusion in this analysis, and will focus on the traditional predictive regression setting. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the unemployment rate in forecasting inflation. The predictive form is of particular interest among policymakers in leveraging the relationship as a practical tool.
@@
Table 8: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ across sample periods and specifications in inflation prediction
(a) TCODE Transformed Data
Sample Period IVX Dlasso XDlasso VR Test
Full Sample -0.014 0.018*** -0.024 0.000
(Jan. 1960 - Apr. 2025) (0.018) (0.006) (0.077)
Pre-Volcker 0.080 0.074*** 0.013 0.125
(Jan. 1960 - Jul. 1979) (0.069) (0.017) (0.224)
Volcker-Greenspan 0.036 -0.020 0.161 0.025
(Aug. 1979 - Jan. 2006) (0.063) (0.020) (0.118)
Bernanke/Yellen/Powell -0.043 -0.002 -0.054 0.002
(Feb. 2006 - Apr. 2025) (0.027) (0.011) (0.120)
(b) Untransformed Data: Excluding I(2) Variables Based on TCODE
Include Labor Variables Exclude Labor Variables
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.077 0.068 0.080 -0.050*** 0.033 0.065
(Jan. 1960 - Apr. 2025) (0.069) (0.218) (0.015) (0.032)
Pre-Volcker -0.129 -0.014 0.522 0.007 0.113 0.548
(Jan. 1960 - Jul. 1979) (0.117) (0.276) (0.051) (0.098)
Volcker-Greenspan 0.094 -0.272 0.741 -0.092* -0.259 0.339
(Aug. 1979 - Jan. 2006) (0.217) (0.287) (0.054) (0.228)
Bernanke-Yellen-Powell 0.550 0.585 0.283 0.001 0.040 0.230
(Feb. 2006 - Apr. 2025) (0.442) (0.975) (0.055) (0.073)
@@
Notes: We report estimates and the standard error (in parentheses below the estimates) across methods and setups. The upper and lower panels corresponds to the scenarios where we use TCODE transformed or untransformed time series as covariates, respectively. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. “VR Test” represents the $p$-value of the variance ratio test (kim2009automatic) on the LASSO residual. The tuning parameter for LASSO estimation is selected through 10-fold block cross-validation. In XDlasso, instruments are generated based on (2.9) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
@@
Table 8 reports the point estimates and standard errors for $\theta_{1}^{*}$ using IVX, Dlasso, and XDlasso across sample periods and specifications. In the benchmark setup with the transformed data, the diagnostic check rejects the m.d.s. condition except for the Pre-Volcker period. Nevertheless, Dlasso, which ignores nonstationarity, delivers significantly positive coefficients for both the full sample and the Pre-Volcker period, in striking contrast to XDlasso and IVX. Using untransformed data significantly alleviates concerns about model misspecification, with $p$-values of the diagnostic test greater than 5% in all cases. Both XDlasso and IVX consistently find no significant predictive power of unemployment for inflation across all specifications and time periods. The empirical findings add new insight to the recent debates on the Phillips curve and echo Mankiw2024’s latest pessimistic remark: “The large confidence intervals for the natural rate, together with the apparent futility of this Holy Grail search, lead me to think that we should not expect much from the Phillips curve as a guide for forecasting inflation or for judging the stance of policy.” On the other hand, with untransformed data, Dlasso yields significantly negative coefficients at the 1% level for the full sample and at the 10% level in the Volcker-Greenspan period, without controlling for other labor market variables. The unstable results of Dlasso across setups highlight the necessity of accounting for nonstationarity in the inference.
@@ Appendix A Technical Proofs
In the proofs, we use $c$ and $C$, without superscripts or subscripts, to denote generic positive constants that may vary place to place. For any positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, $a_{n}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}b_{n}$” means that there is an absolute constant, say $c$, such that the event $\left\{a_{n}\leq cb_{n}\right\}$ holds with probability approaching one (w.p.a.1.). Symmetrically, $a_{n}\stackrel{{\scriptstyle\mathrm{p}}}{{\succcurlyeq}}b_{n}$” means $b_{n}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}a_{n}$”. The integer floor function is denoted as $\left\lfloor\cdot\right\rfloor$. For an $n$-dimensional vector $x=(x_{t})_{t\in[n]}$, the $L_{2}$-norm is $\left\|x\right\|_{2}=\sqrt{\sum_{t=1}^{n}x_{t}^{2}}$. For notational simplicity, in the proofs we assume $p\geq n^{\nu_{1}}$ for some absolute constant $\nu_{1}$, which is reasonable as we focus on the high dimensional case with a larger $p$ relative to $n$.A.1A.1A.1There is no technical difficulty in allowing $p$ to grow either slowly at a logarithmic or fast at an exponential rate of $n$, but we have to compare $\log p$ and $\log n$ in many places, and in many conditions and rates the term $\log p$” has to be changed into $\log(np)$.We use $\mathbf{I}_{n}$ to denote the $n$-dimensional identity matrix, where the subscript may be omitted when there is no ambiguity with matrix dimensions.
Section A.1 proves the results in Section 3.1 for the consistency of Slasso at the presence of both LUR and stationary regressors. Section A.2 includes the proofs for the results in Section 3.2 that constructs consistency of the auxiliary LASSO regression adapted for bias correction. Section A.3 includes the proofs for the theorems in Section 3.3 about the asymptotic distribution of the XDlasso estimator and the order of its standard error, which determine the size and power of the XDlasso inference.
A.1 Proofs for Section 3.1
A.1.1 Technical Lemmas of Gaussian Approximation
The assumptions imposed in Section 3 slightly differ from those in mei2022lasso (MS24, hereafter). Specifically, the linear process assumption in MS24 is extended to more general $\alpha$-mixing and sub-exponential conditions. One of the main modifications of the proof is the Gaussian approximation error deduced in MS24’s Lemma B.4. The following lemma re-establishes Gaussian approximation errors for the stationary components $v_{t}$ defined as (3.1). The results of Gaussian approximation will be useful for the RE condition required for the Slasso’s consistency.
@@ A.1.1 Technical Lemmas of Gaussian Approximation
$\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(t\cdot V_{k}^{*})\right)\right|=O_{p}\left(\dfrac{(\log p)^{3/2}}{n^{1/4}}\right)$
@@
The convergence rate in Lemma A.1 is less sharp than that in MS24’s Lemma B.4, since we work with general $\alpha$-mixing sequences without specifying linear processes.
@@
Recall that Assumption 2 imposes an upper bound for the $\alpha$-mixing coefficient. With $p=q=3$ in Equation (2.2) in the Corollary of davydov1968convergence, we have
@@
Step 2. Verifying (A.2). We use the strong Gaussian approximation from lin1997limit’s Theorem 9.3.1. Specifically, define $g(x)=\exp(x)$. By the sub-exponential tail imposed by Assumption 1, the sub-exponential norm of $\varepsilon_{k,t}$, denoted as $\|\varepsilon_{k,t}\|_{g}$ in lin1997limit, is uniformly bounded by an absolute constant. It then suffices to verify the following two conditions required in the aforementioned theorem:
@@
    $\sum_{d=1}^{\infty}\alpha(d)^{1/4}\cdot\log(1/\alpha(d))<\infty$, where the parameter $\delta$ in lin1997limit is taken as 2.
Then by Theorem 9.3.1 of lin1997limit, for any $j\in[p]$
@@
Proof of (i). By (A.4) and (A.5),
$\lim_{t\to\infty}\sup_{k\in[p]}\left|\dfrac{V_{k,t}}{t}-V_{k}^{*}\right|$ $\leq\lim_{t\to\infty}2\sup_{k\in[p]}\left|\sum_{d=t+1}^{\infty}{\rm cov}_{k}(d)+\dfrac{2}{t}\sum_{d=1}^{t}d\cdot{\rm cov}_{k}(d)\right|$
$\leq\lim_{t\to\infty}C\left[\sum_{d=t+1}^{\infty}\exp\left(-c_{\alpha}d^{r}/2\right)+\dfrac{2}{t}\sum_{d=1}^{t}d\cdot{\rm cov}_{k}(d)\right]$
$=0,$
where $C$ is an absolute constant, and the limit applies (A.6) and (A.7). By Assumption 2, the long-run variance $V_{k}^{*}$ is bounded from below by some absolute constant. This result implies that $V_{k,t}/t$ is lower bounded by some absolute constant uniformly for all $(k,t)$.